UkuleleGirl:

Idk what this is. Please help :3

UkuleleGirl:

Cyber:

@tetsxpreme Please help this user

jacob65:

Alright i can help if your ok witht hat

UkuleleGirl:

I don't think tet knows this.

Cyber:

I think a mathlete should know

UkuleleGirl:

ok

jacob65:

I can help

UkuleleGirl:

please do.

jacob65:

Alright are we solving for x?

UkuleleGirl:

I think were supposed to factor.

UkuleleGirl:

I need dude online. He knows all about this.

jacob65:

ok go pm him then

UkuleleGirl:

Hes not on.

jacob65:

Oh alright

UkuleleGirl:

oh wait he is.

UkuleleGirl:

@dude

UkuleleGirl:

-_- he went offline.

jacob65:

lol i dont remember how to fractor it its been years but I bet i could help solve it

UkuleleGirl:

I know how to factor

jacob65:

then whats the issue

UkuleleGirl:

I always get stuck on here. Don't question me.

jacob65:

bruh dont ask for help then lol

UkuleleGirl:

I need help....breh. You don't get to tell me what to do ON MY question.

UkuleleGirl:

Get out if your not going to help

YTStresssPG3D:

@ukulelegirl wrote:
Get out if your not going to help
they said they was going to help u

YTStresssPG3D:

like chill

UkuleleGirl:

They were being rude and im not in the mood. Just help me please...

UkuleleGirl:

I dont' feel like gettin yelled at again by my mom

YRJ8498:

@ukulelegirl wrote:
I dont' feel like gettin yelled at again by my mom
LMAO

YTStresssPG3D:

@jacob65 wrote:
Alright i can help if your ok witht hat
uh no because of that

UkuleleGirl:

THEN LEAVE IM NOT IN THE MOOD

UkuleleGirl:

Just go...

YTStresssPG3D:

shoula let thm help u

UkuleleGirl:

OK just go. This convo is over.

Cyber:

Here I will help you get someone @ukulelegirl

Cyber:

@joe348

UkuleleGirl:

Your supposed to factor-

YTStresssPG3D:

@ukulelegirl wrote:
Your supposed to factor-
no direct answers

UkuleleGirl:

but this always confusesssss meeeeeeeee T_T

UkuleleGirl:

me and you both know thats not the answer

RodWav3:

oh damn

ThisGirlPretty:

For those that don't know how to help the user please refrain from making any comments. Instead it would be more helpful if we pinged the users that were good in math.

UkuleleGirl:

Exactly.

UkuleleGirl:

This is why we need dude. or someone smart like tet, joe, noodles, them peoople

jacob65:

I offered to help and she continued to be rude about it so i left

Joe348:

Well have you tried solving the equation? @ukulelegirl

YTStresssPG3D:

@jacob65 wrote:
I offered to help and she continued to be rude about it so i left
yes he was trying to answer but she was being rude

UkuleleGirl:

ok if your going to start on me jacob i can be rude. now leave. thanks. im trying to learn.

YTStresssPG3D:

or she

UkuleleGirl:

And no joe...I'm confused...i would say combine like terms but there isn't anything u can combin here

YTStresssPG3D:

@ukulelegirl wrote:
ok if your going to start on me jacob i can be rude. now leave. thanks. im trying to learn.
he was trying to help u if u dont want help from people offering it then take the question down.

jacob65:

Im a she lol and i did leave and ur still going i tried to help yet you were being the rude one seriosuly

YTStresssPG3D:

@jacob65 wrote:
Im a she lol and i did leave and ur still going i tried to help yet you were being the rude one seriosuly
i said or she and yes

iFish:

I can answer this question, if you guys would quit talking.

jacob65:

thank you

UkuleleGirl:

JUST SHUT UP UNLESS U CAN HELP DANG

YTStresssPG3D:

@ifish wrote:
I can answer this question, if you guys would quit talking.
she doesnt want help from anyone thats not smart she will not take your answer

UkuleleGirl:

its not that.

Joe348:

@ukulelegirl wrote:
And no joe...I'm confused...i would say combine like terms but there isn't anything u can combin here
your not combining your trying to solve for the equation first

TETSXPREME:

@hero

UkuleleGirl:

I know were not. which is why im confused. Im stuck.

iFish:

I can answer this question, just let me speak.

UkuleleGirl:

go ahead

iFish:

Do you know how to factor the equation?

UkuleleGirl:

It says this is factor by grouping.

iFish:

Grab a sheet of paper, and write this information down, please.

UkuleleGirl:

ok brb

UkuleleGirl:

alr im here.

UkuleleGirl:

@yrj8498 wrote:
@ytstressspg3d wrote:
@ifish wrote:
I can answer this question, if you guys would quit talking.
she doesnt want help from anyone thats not smart she will not take your answer
Stress leave the question so she could get a real answer and not ppl calling her rude I would want sm1 smart answering my question too
Exactly thank you

Joe348:

\[(x+1)(2x+1)(2x-1)=0\]

YTStresssPG3D:

@ukulelegirl wrote:
@yrj8498 wrote:
@ytstressspg3d wrote:
@ifish wrote:
I can answer this question, if you guys would quit talking.
she doesnt want help from anyone thats not smart she will not take your answer
Stress leave the question so she could get a real answer and not ppl calling her rude I would want sm1 smart answering my question too
Exactly thank you
well someone was just trying to help and she continued to be rude.

UkuleleGirl:

you factored out 2 from 4?

ThisGirlPretty:

Guys let's stay on topic please

UkuleleGirl:

Thank you tgp. <3

UkuleleGirl:

@joe348 wrote:
\[(x+1)(2x+1)(2x-1)=0\]
DId you factor out 2 from 4

UkuleleGirl:

@hero

UkuleleGirl:

Joe is that the answer or???

UkuleleGirl:

I'm confused..

jacob65:

bro be patient

Joe348:

sorry, im a lil laggy

UkuleleGirl:

I'm sorry this is making me very frustrated...Please excuse me If I go offline for a bit. I just am about to lose it.

UkuleleGirl:

and your fine joe

Hero:

Someone posted the answer but they did not provide any steps or explanations. Really sad.

UkuleleGirl:

Yes.

UkuleleGirl:

What sucks is i have more like these.

ThisGirlPretty:

Think you can help her hero?

YRJ8498:

@tetsxpreme

Timmyspu:

Could you provide the steps for why that is the answer hero?

Hero:

Yes

CheeseStick7793:

i c an helppppppp

UkuleleGirl:

Knock yourself out.

Hero:

So we have \(4x^3 + 4x^2 - x - 1 = 0\) and we have to factor by grouping. So the first step is to factor the first two terms and the last two terms. \(4x^2\) can be factored from the first two terms and \(-1\) can be factored from the last two terms. Upon doing so you get: \(4x^2(x + 1) -1(x + 1)\) Next, observe that \(x+1\) is common to both factorizations, so factor that out as well to get: \((x+1)(4x^2 -1)\) Finally, observe that \(4x^2 - 1\) is a difference of squares. Then factor it to get: \((x+1)(2x+1)(2x-1)\)

Joe348:

@joe348 wrote:
\[(x+1)(2x+1)(2x-1)=0\]
\[4x^3+x^2)+(-4x-1)=0\] \[x^2(4x+1)-1(4x+1)=0\] \[(x^2-1)(4x+1)=0\]

Joe348:

@joe348 wrote:
@joe348 wrote:
\[(x+1)(2x+1)(2x-1)=0\]
\[4x^3+x^2)+(-4x-1)=0\] \[x^2(4x+1)-1(4x+1)=0\] \[(x^2-1)(4x+1)=0\]
(4x^3)

CheeseStick7793:

(x+1)(2x-1)(2x+1)

UkuleleGirl:

uhhhhhhh.

CheeseStick7793:

trust me

Joe348:

That's how I got it but it's not the answer

CheeseStick7793:

that is your answer

CheeseStick7793:

(x+1)(2x-1)(2x+1)

UkuleleGirl:

OMG. its so confusing.

CheeseStick7793:

mine is right

UkuleleGirl:

I think I'm really going to have to do more of these to understand a bit more

CheeseStick7793:

Answer: (x+1)(2x-1)(2x+1)

Hero:

The steps

@hero wrote:
So we have \(4x^3 + 4x^2 - x - 1 = 0\) and we have to factor by grouping. So the first step is to factor the first two terms and the last two terms. \(4x^2\) can be factored from the first two terms and \(-1\) can be factored from the last two terms. Upon doing so you get: \(4x^2(x + 1) -1(x + 1)\) Next, observe that \(x+1\) is common to both factorizations, so factor that out as well to get: \((x+1)(4x^2 -1)\) Finally, observe that \(4x^2 - 1\) is a difference of squares. Then factor it to get: \((x+1)(2x+1)(2x-1)\)
@UkuleleGirl, the steps I provided are the correct steps.

CheeseStick7793:

you are wrong

UkuleleGirl:

How is he wrong? hes literally like one of the smartest people here.

CheeseStick7793:

beacuse i calculated and checked my work

CheeseStick7793:

and i got (x+1)(2x-1)(2x+1)

UkuleleGirl:

Thats what he got.

CheeseStick7793:

he did not g et the same the math symbols are switched

Hero:

BTW, the problem is still not finished @UkuleleGirl. You still have to solve for \(x\)

UkuleleGirl:

Set them equal to 0 and solve yes i know this part

CheeseStick7793:

(x+1)(2x-1)(2x+1)

chrisdjackson:

U still need help with this equation?

UkuleleGirl:

is the answers, \[-1, -\frac{ 1 }{ 2 },\frac{ 1 }{ 2 }\]

Hero:

Yes correct @UkuleleGirl

UkuleleGirl:

Yay!

UkuleleGirl:

Smh math is hard

CheeseStick7793:

Factor by grouping Factor by grouping Factor by grouping Factor by grouping (x+1)(4x^2-1) Find one factor Factor by grouping Factor by grouping Factor by grouping Factor by grouping (x+1)(2x-1)(2x+1)